# Catch Me if You Can: Detecting Unauthorized Data Use in Deep Learning Models

### v2

@@ Abstract.
MembershipTracker only requires the users to mark a small fraction of data (0.005$\sim$0.1in proportion to the training set), and it enables the users to reliably detect the unauthorized use of their data (average 0% FPR@100% TPR). We show that MembershipTracker is highly effective across various settings, including industry-scale training on the full-size ImageNet-1k dataset. We finally evaluate MembershipTracker under multiple classes of countermeasures.
@@ 1. Introduction
  • Develop a high-power set-based MI verification process that can reliably audit whether the target data are used to train the model (with low FPR and high TPR).
@@
We find that by merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 0% FPR@100% TPR), and it is scalable to the large-scale ImageNet training (on both Convolutional and Transformer network). To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting. This renders MembershipTracker a practical data auditing tool that contributes to responsible AI practice.
@@ 2. Related Work
Wenger et al. propose to inject a target spurious feature into the user data to be learned by the suspected model, and then detect whether the model has associated the injected feature with the target class label (Wenger et al., 2024). To detect, the users can separately overlay the target and non-target features to some auxiliary data (outside the target class), and then determine whether the target feature causes a slightly higher prediction probability on the target class, compared with the non-target features (e.g., 10% vs. 1%). Both their work and ours involve adding spurious features to mark users’ data, but there are several major differences between the two. Notably, our analysis reveals that they (Wenger et al., 2024) significantly underestimate their detection false positives (from 0% to >30%). This is because they employ a discrepant detection procedure for the testing features that are used for training (true positives), and those that are not (false positives); however, a fair evaluation should follow a consistent procedure (see our detailed analysis in Appendix D).
@@
Instance-level MI detects whether a specific instance is used to train a model. It was first proposed by Shokri et al. (Shokri et al., 2017) and there are many follow-up studies (Yeom et al., 2018; Choquette-Choo et al., 2021; Bertran et al., 2024; Ye et al., 2022; Carlini et al., 2022). However, even state-of-the-art MI methods (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022) still struggle with achieving high TPR when controlled at the low FPR regime (more in Section 4.1).
This has led to several attempts at improving the effectiveness of existing MI methods via data poisoning. Tramer et al. propose to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence on the model’s decision, thereby making the target samples easier to be de-identified (Tramèr et al., 2022). While injecting mislabeled samples represents an effective solution for improving the MI success (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023), it may be challenging for the users to apply to their data, because in many scenarios the users may not have control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that such a method can still only increase the MI success to a limited extent (validated in Appendix E). By contrast, MembershipTracker does not assume control of the data labels, and is still able to facilitate high-power MI for tracing data provenance.
@@
Recent work by Huang et al. (Huang et al., 2024) proposes a general framework for auditing unauthorized data use in training DL models. Their idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published vs. the unpublished one. While their framework can work for different domains such as foundation models, they still require the data holders to mark a substantial portion (1$\sim$10%) of dataset. Instead, MembershipTracker considers the more challenging (and realistic) setting where the marked data amounts to $\leq 0.1\%$ of the dataset. We also “stress test” their technique under a similar setting as ours and find that their performance degrades considerably then (see Appendix F).
@@ 3.1.1. User capability
Next, in the provenance verification process, we assume $U$ has access to a set of non-member samples that are not used for training the target model [^1] Under the presumption that only a small fraction of data sampled from the distribution were used in training, one may simply draw a random sample from the underlying distribution, and be confident that it is representative of the non-member data (Bertran et al., 2024; Ye et al., 2022)., which is a common assumption [^2] While we also adopt this common assumption in our study, we conduct additional experiments and find that this assumption can in fact, be relaxed (see Section 7). to control the MI process at a low FPR regime (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022). $U$ also has black-box access to query the target model. In a real world setting, users often have limited control and knowledge beyond their own data, and thus we specify the constraints by $U$ as follows.
@@
3.1.2. Evaluation metric
We follow the best practice to measure the MI success by its true positive rate (TPR) and false positive rate (FPR) (Carlini et al., 2022), which is fully characterized by the Receiver Operating Characteristic (ROC) curve. From the ROC curve, we can compute either 1g11g TPR under a fixed FPR; or 1g21g FPR under a fixed TPR. We use metric 1g21g, and we first explain why 1g11g is unsuitable for evaluating the data-provenance based MI method in our work.
Motivation. The goal of performance evaluation is to understand whether MembershipTracker can achieve both low FPR (to avoid false accusation) and high TPR (for correct data provenance). However, manually choosing a low FPR threshold and evaluating the corresponding TPR (as in 1g11g) can be undesirable, as the pre-defined FPR threshold may be too conservative or aggressive. We use Fig. 2 (from our experiment) to explain the tradeoffs involved.
For the left method in Fig. 2, using a 1% FPR threshold is too conservative. Under this threshold, both methods in Fig. 2 will be deemed to have the same performance, yet the left method is in fact much more powerful and it can yield the same 100% TPR under a lower FPR of 0% (undesirable).
Likewise, setting a tighter FPR threshold such as 0%, is appropriate for the left method, but this becomes too aggressive for the right method in Fig. 2 - this high-power method (with 0.9993 AUC) yields a poor outcome of 5% TPR (undesirable).
Figure 2. From the same ROC curve, we can compute either 1g11g TPR@fixed FPR; or 1g21g FPR@fixed TPR.
In 1g11g, the pre-defined FPR threshold can be too conservative (✗ on the left) or too aggressive (✗ on the right).
Instead, metric 1g21g is able to properly reflect the high performance by both methods (with low FPRs and high TPRs), and thus we adopt it for performance evaluation purposes.
Proposed metric. To overcome the above issue, we advocate the use of metric 1g21g (FPR under a fixed TPR).
Under a 100% TPR, the left method in Fig. 2 has a 0% FPR, and the right method has a slightly higher 0.52% FPR. This result shows that (1) both methods have very good performance (in having low FPRs and high TPRs); and (2) the left method is slightly better than the right one (0% vs 0.52% FPR).
Based on the above, we adopt FPR@fixed TPR as the evaluation metric, and we set a 100% TPR threshold to simulate the case where all the users whose data are misused without their consent can trace the provenance of their data (see our design goal in Section 4). An ideal technique should incur as low FPR as possible.
Remark. Note that this metric serves to measure the minimal FPR for MembershipTracker to achieve 100% TPR. We will show later that MembershipTracker has very low FPR@100% TPR. Therefore, this result represents a desirable FPR threshold that can be used to calibrate the actual MI data auditing process, which follows the common practice to control the MI process under the low FPR regime, and detects whether the testing data was used for training or not (details in Section 4.4.3).
For instance, if MembershipTracker achieves 0.1% FPR@100% TPR, it means users can control the MI process under 0.1% FPR [^3] Note that users can still choose an arbitrary FPR threshold (e.g., 0% or 1% FPR)., and MembershipTracker will be able to detect those data used for training (with 100% TPR).
In essence, metric 1g11g and metric 1g21g merely represent specific points in the ROC curve, and a full performance characterization should still be captured by the ROC curve. However, for MembershipTracker, we find that using metric 1g21g is already able to illustrate that MembershipTracker can achieve high performance in the low FPR regime. For example, our evaluation in Section 5.1 shows that MembershipTracker achieves an average of 0% FPR@100% TPR across different evaluation settings, which means MembershipTracker can achieve perfect discrimination in the low FPR regime. In these cases, all ROC curves will show a flat line similar to that on the left of Fig. 2, and thus we omit the curves for space reasons.
With that said, in evaluating other related techniques (Carlini et al., 2022; Wenger et al., 2024; Tramèr et al., 2022), we have carefully provided the full ROC curve for a comprehensive comparison (e.g., Fig. 3, Fig. 15, and Fig. 17).
@@ 4. Methodology
We first describe our design goals, and then conduct an empirical study to understand why directly leveraging existing MI methods cannot meet our goals. We then present MembershipTracker with a data marking and provenance verification process: first by marking the users’ protected data with targeted changes (Section 4.3); then performing a specialized MI process to audit whether the marked data are used to train the model (Section 4.4).
@@
Goal (1): High provenance effectiveness. Our main goal is to enable those users whose data are misused for model training to perform successful data provenance (100% TPR), and we also want to ensure low FPR to avoid false accusation.
@@
4.1. Motivation
Our work addresses data provenance through MI, and there are several available MI methods (Ye et al., 2022; Bertran et al., 2024; Carlini et al., 2022; Tramèr et al., 2022). However, these methods still suffer from limited MI effectiveness and/or impose strong assumptions on the users - we elaborate both limitations next.
1g11g
Existing MI methods have limited provenance effectiveness, and incur very high FPR under 100% TPR.
To validate this, we use two methods in the state-of-the-art Likelihood-ratio attack, one is with and the other without shadow model calibration (128 shadow models) (Carlini et al., 2022). We evaluate both methods on a WideResNet-28-4 model trained on half of the training set in CIFAR100 dataset, and they both incur a high FPR of 75.32% and 64.62%, under 100% TPR.
With that said, the above results should not be interpreted as the failure of these approaches, because they were originally built from an adversarial perspective to expose the worst-case privacy leakage in DL models, where even just reliably de-identifying a few member samples is considered meaningful (Carlini et al., 2022), e.g., the shadow-model based method in Fig. 3 can achieve 27.21% TPR@0.1% FPR.
On the other hand, our work considers MI for data provenance purposes, which represents a more challenging task, and requires both low FPR and high TPR. The high TPR is crucial for the users to be able to detect the unauthorized use of their data, while low FPR is important to minimize the chance of false accusation.
As in Fig. 3, existing MI methods struggle with this task. The reason is that the training samples are not always strongly memorized by the model (except some outliers) (Yeom et al., 2018; Carlini et al., 2022; Tramèr et al., 2022), which is directly related to the ability of performing accurate MI. This problem motivates related work (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023) to amplify the model’s memorization on the training samples, and we discuss their limitation next.
Figure 3. MI for tracing data provenance requires both high TPR (for correct provenance detection), and low FPR (to avoid false accusation). However, even state-of-the-art MI methods (Carlini et al., 2022) cannot meet this challenging goal, and are unsuitable for data provenance purposes.
1g21g
There are other methods that do not require shadow-model calibration, but they either suffer from poor effectiveness (Yeom et al., 2018; Carlini et al., 2022) or impose unrealistic assumption on the users (Tramèr et al., 2022; Chaudhari et al., 2023). E.g., the approach by Tramer et al. (Tramèr et al., 2022) can improve the MI success without shadow-model calibration, but it assumes users have the ability to mislabel some training data, which can be challenging in the cases where the users have no control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that the increase of MI success is still largely limited (see Appendix E for a validation).
4.2. Overview
We next present MembershipTracker, a technique that requires neither shadow-model training nor manipulating data labels, and can still achieve high-power MI (with low FPR and high TPR) for tracing data provenance. It consists of a data marking and a MI-based verification process - we explain them next.
@@ 4.3.1. An initial strawman approach
There are various ways to generate OOD samples, and we develop a simple method that creates samples with random color stripes, such as the one shown below (the method’s details are in Section 5). Meanwhile, other alternative methods such as using samples from an OOD dataset may also be considered (Appendix C.1.3).
@@
To perform MI, we compare the prediction loss on the target OOD samples (from the member users) and some non-target OOD samples (from the non-member users). We find that the model indeed strongly memorizes those OOD samples present in the training set, for which we observe a 0% FPR@100% TPR. In comparison, if the target data are unmodified, the model only exhibits weak memorization on them, which yields 56.32% FPR@100% TPR.
@@ 4.3.2. Step 1: Image blending
As discussed earlier, the marked samples should contain the original feature (to preserve visual information), as well as the OOD feature (for provenance purposes). Inspired by the common image blending technique (Zhang et al., 2018; Wenger et al., 2024), we propose to blend the target sample with the OOD feature:
@@
where $m$ moderates the contribution of different features.
By using a large $m$, we can largely preserve the high image quality (e.g., the center image in Fig. 4 is marked with $m=0.7$) while keeping the OOD feature to amplify the model’s memorization on the resulting samples. We find that merely blending the OOD feature into the target samples ($m=0.7$) can reduce the FPR@100% TPR from 56.32% to 23.58%.
Limitation. Although the FPR is largely reduced, it is still too high. We find that this is due to the existence of the original feature, which hinders the model’s memorization on the OOD feature. We next explain how to mitigate this.
@@ 4.3.3. Step 2: Noise injection
Generating the perlin noise is a lightweight process that does not assume access to a proxy model. Despite its simplicity, we find that injecting a small amount of perlin noise (with $\delta=8/255$) is very effective and it can further decrease the FPR from 23.58% to 7.97% (a 66% reduction).
Overall, the two-step marking process is able to reduce the FPR from 56.32% to 7.97%. While the FPR can be further reduced by using a smaller $m$ for image blending or larger $\delta$ for noise injection, it would increase the visual distortion (see Appendix C.1.2 for an evaluation). We thus propose another innovation in the MI process to mitigate the FPR, and without degrading the image quality.
@@ 4.4. Data Provenance via Membership Inference
To start with, we note that so far we have been considering MI on a common instance basis (Shokri et al., 2017; Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022), which uses the per-instance loss as the signal function value for MI. We first explain why this approach would incur high FPR, and then present a specialized MI process to overcome it.
4.4.1. Why instance-based MI suffers from high FPR
It uses the per-sample loss for MI, which can incur high FPR even if the prediction loss on the target marked samples are very low. This is because there may exist many non-member samples with low loss values as well, which would be mis-identified as member samples, and result in high FPR.
For instance, the left figure in Fig. 5 shows the individual loss values on different samples, where many non-member samples also yield low prediction loss (the low loss values are shown towards the right side because they are plotted in the logit-scaled form (Carlini et al., 2022) for visualization purposes). This leads to a 7.97% FPR @100% TPR.
Figure 5. Comparison of using the per-instance loss and the average loss across the samples by each user (each contributes 0.1% of the training data) as the signal function for MI verification. The former yields 7.97% FPR@100% TPR; while the latter has 0% FPR@100% TPR (our proposal).

### v3

# Anonymity Unveiled: A Practical Framework for Auditing Data Use in Deep Learning Models

@@ Abstract.
MembershipTracker only requires the users to mark a small fraction of data (0.005$\sim$0.1in proportion to the training set), and it enables the users to reliably detect the unauthorized use of their data (average 100% TPR@0% FPR). We show that MembershipTracker is highly effective across various settings, including industry-scale training on the full-size ImageNet-1k dataset. We finally evaluate MembershipTracker under multiple classes of countermeasures [^1] Our code is available at https://github.com/DependableSystemsLab/MembershipTracker..
Data auditing, data provenance, membership inference
journalyear: 2025copyright: acmlicensedconference: ACM Conference on Computer and Communications Security; October 13–17, 2025; Taipei, Taiwanbooktitle: ACM Conference on Computer and Communications Security (CCS’25), October 13–17, 2025, Taipei, Taiwanccs: Security and privacyccs: Computing methodologies Machine learning
@@ 1. Introduction
  • Develop a high-power set-based MI verification process that can reliably audit whether the target data are used to train the model (with high TPR when controlled under a low FPR).
@@
We find that by merely marking a small fraction of samples (0.005$\sim$0.1of the dataset), MembershipTracker effectively empowers the users to reliably audit the usage of their data (average 100% TPR@0% FPR), and it is scalable to the large-scale ImageNet training (on both Convolutional and Transformer network). To the best of our knowledge, MembershipTracker is the first technique that can scale to support reliable data auditing in large-scale training setting. This renders MembershipTracker a practical data auditing tool that contributes to responsible AI practice.
@@ 2. Related Work
Wenger et al. propose to inject a target spurious feature into the user data to be learned by the suspected model, and then detect whether the model has associated the injected feature with the target class label (Wenger et al., 2024). To detect, the users can separately overlay the target and non-target features to some auxiliary data (outside the target class), and then determine whether the target feature causes a slightly higher prediction probability on the target class, compared with the non-target features (e.g., 10% vs. 1%). Both their work and ours involve adding spurious features to mark users’ data, but there are several major differences between the two. Notably, our analysis reveals that they (Wenger et al., 2024) significantly underestimate their detection false positives (from 0% to >30%). This is because they employ a discrepant detection procedure for the testing features that are used for training (true positives), and those that are not (false positives); however, a fair evaluation should follow a consistent procedure (see our detailed analysis in Appendix C).
@@
Instance-level MI detects whether a specific instance is used to train a model. It was first proposed by Shokri et al. (Shokri et al., 2017) and there are many follow-up studies (Yeom et al., 2018; Choquette-Choo et al., 2021; Bertran et al., 2024; Ye et al., 2022; Carlini et al., 2022). However, even state-of-the-art MI methods (Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022) still struggle with achieving high TPR when controlled at the low FPR regime (more in Section LABEL:sec:motivation).
This has led to several attempts at improving the effectiveness of existing MI methods via data poisoning. Tramer et al. propose to inject mislabeled samples into the model’s training set, which can transform the target samples into outliers and amplify their influence on the model’s decision, thereby making the target samples easier to be de-identified (Tramèr et al., 2022). While injecting mislabeled samples represents an effective solution for improving the MI success (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023), it may be challenging for the users to apply to their data, because in many scenarios the users may not have control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that such a method can still only increase the MI success to a limited extent (validated in Appendix D). By contrast, MembershipTracker does not assume control of the data labels, and is still able to facilitate high-power MI for tracing data provenance.
@@
Recent work by Huang et al. (Huang et al., 2024) proposes a general framework for auditing unauthorized data use in training DL models. Their idea is to generate two perturbed copies of the target data, randomly publish one of them, and then compare the model’s membership score on the published vs. the unpublished one. While their framework can work for different domains such as foundation models, they still require the data holders to mark a substantial portion (1$\sim$10%) of dataset. Instead, MembershipTracker considers the more challenging (and realistic) setting where the marked data amounts to $\leq 0.1\%$ of the dataset. We also “stress test” their technique under a similar setting as ours and find that their performance degrades considerably then (see Appendix E).
@@ 3.1. MI for Data Provenance
Unless otherwise stated, we follow the common practice to measure the MI success by its true positive rate (TPR) controlled under a low false positive rate (FPR) (Carlini et al., 2022; Ye et al., 2022).
@@ 3.1.1. User capability
Next, in the provenance verification process, we assume $U$ has access to a set of non-member samples that are not used for training the target model [^2] Under the presumption that only a small fraction of data sampled from the distribution were used in training, one may simply draw a random sample from the underlying distribution, and be confident that it is representative of the non-member data (Bertran et al., 2024; Ye et al., 2022)., which is a common assumption [^3] While we also adopt this common assumption in our study, we conduct additional experiments and find that this assumption can in fact, be relaxed (see Section 6.2). to control the MI process at a low FPR regime. $U$ also has black-box access to query the target model. In a real world setting, users often have limited control and knowledge beyond their own data, and thus we specify the constraints by $U$ as follows.
@@ 4. Methodology
We first conduct an empirical study to understand the limitations of leveraging existing MI methods for data auditing, and then describe our design goals. We then present MembershipTracker with a data marking and provenance verification process: first by marking the users’ protected data with targeted changes (Section 4.1); then performing a specialized MI process to audit whether the marked data are used to train the model (Section 4.2).
Motivation. Our work addresses data provenance through MI, and there are several available MI methods (Ye et al., 2022; Bertran et al., 2024; Carlini et al., 2022; Tramèr et al., 2022). However, these methods still suffer from limited MI effectiveness in the low FPR regime and/or impose strong assumptions on the users - we elaborate both limitations next.
1g11g
Existing MI methods achieve limited TPR under a low FPR.
To validate this, we use two methods in the state-of-the-art Likelihood-ratio attack, one is with and the other without shadow model calibration (128 shadow models) (Carlini et al., 2022). We evaluate both methods on a WideResNet-28-4 model trained on half of the training set in CIFAR100 dataset. As shown, even the shadow-model based method can only achieve 27.21% TPR@0.1% FPR.
This is undesirable because achiving high TPR (under a low FPR) is crucial for the users to reliably detect the unauthorized use of their data. As in Fig. 2, existing MI methods struggle with this task. The reason is that the training samples are not always strongly memorized by the model (except some outliers) (Yeom et al., 2018; Carlini et al., 2022; Tramèr et al., 2022), which is directly related to the ability of performing accurate MI. This problem motivates related work (Tramèr et al., 2022; Chen et al., 2022; Chaudhari et al., 2023) to amplify the model’s memorization on the training samples, and we discuss their limitation next.
Figure 2. State-of-the-art MI methods (Carlini et al., 2022) achieve limited TPR under the low FPR regime (undesirable).
1g21g
There are other methods that do not require shadow-model calibration, but they either suffer from poor effectiveness (Yeom et al., 2018; Carlini et al., 2022) or impose unrealistic assumption on the users (Tramèr et al., 2022; Chaudhari et al., 2023). E.g., the approach by Tramer et al. (Tramèr et al., 2022) can improve the MI success without shadow-model calibration, but it assumes users have the ability to mislabel some training data, which can be challenging in the cases where the users have no control over the data labels (Dzieza, 2023; Wenger et al., 2024; Deng et al., 2009). Even if this is feasible, we find that the increase of MI success is still largely limited (see Appendix D for a validation).
@@
Goal (1): High provenance effectiveness. Our main goal is to enable those users whose data are misused for model training to perform successful data provenance (high TPR), and we also want to ensure low FPR to avoid false accusation. Therefore, our goal is to achieve high TPR@low FPR.
@@ 4.1.1. An initial strawman approach
There are various ways to generate OOD samples, and we develop a simple method that creates samples with random color stripes, such as the one shown below (the method’s details are in Section 5). Meanwhile, other alternative methods such as using samples from an OOD dataset may also be considered (Appendix B.1.3).
@@
To perform MI, we compare the prediction loss on the target OOD samples (from the member users) and some non-target OOD samples (from the non-member users). We find that the model indeed strongly memorizes those OOD samples present in the training set, for which we observe a 100% TPR@1% FPR. In comparison, if the target data are unmodified, the model only exhibits weak memorization on them, which yields merely 1.28% TPR@1% FPR.
@@ 4.1.2. Step 1: Image blending
As discussed earlier, the marked samples should contain the original feature (to preserve visual information), as well as the OOD feature (for provenance purposes). Inspired by the common image blending technique (Zhang et al., 2018a; Wenger et al., 2024), we propose to blend the target sample with the OOD feature:
@@
where $x$ is the original image, $x_{ood}$ is the OOD feature, and $m$ moderates the contribution of different features.
By using a large $m$, we can largely preserve the high image quality (e.g., the center image in Fig. 3 is marked with $m=0.7$) while keeping the OOD feature to amplify the model’s memorization on the resulting samples. We find that merely blending the OOD feature into the target samples ($m=0.7$) can improve the TPR@1% FPR from 1.28% to 16.44%.
Limitation. Although the TPR is largely improved, it is still too low. We find that this is due to the existence of the original feature, which hinders the model’s memorization on the OOD feature. We next explain how to mitigate this.
@@ 4.1.3. Step 2: Noise injection
Generating the perlin noise is a lightweight process that does not assume access to a proxy model. Despite its simplicity, we find that injecting a small amount of perlin noise (with $\delta=8/255$) is very effective and it can further increase the TPR@1% FPR from 16.44% to 57.8% (a 3.5x improvement).
Overall, the two-step marking process is able to increase the TPR@1% FPR from 1.28% to 57.8%. While the TPR can be further improved by using a smaller $m$ for image blending or larger $\delta$ for noise injection, it would increase the visual distortion (see Appendix B.1.2 for an evaluation). We thus propose another innovation in the MI process to improve the TPR, and without degrading the image quality.
@@ 4.2. Data Provenance via Membership Inference
To start with, we note that so far we have been considering MI on a common instance basis (Shokri et al., 2017; Carlini et al., 2022; Bertran et al., 2024; Ye et al., 2022), which uses the per-instance loss as the signal function value for MI. We first explain why this approach would incur low TPR under a low FPR, and then present a specialized MI process to overcome it.
4.2.1. Why instance-based MI suffers from low TPR
It uses the per-sample loss for MI, which can incur low TPR even if the prediction loss on the target marked samples are very low. This is because many non-member samples may also have low loss values. To control the MI process under a low FPR, only a small fraction of member samples with significantly lower prediction loss are recognized as members, thereby resulting in low TPR.
For instance, the left figure in Fig. 4 shows the individual loss values on different samples, where many non-member samples also yield low prediction loss (the low loss values are shown towards the right side because they are plotted in the logit-scaled form (Carlini et al., 2022) for visualization purposes). This leads to a 57.8% TPR@1% FPR.
Figure 4. Comparison of using the per-instance loss and the average loss across the samples by each user (each contributes 0.1% of the training data) as the signal function for MI verification. The former yields 57.8% TPR@1% FPR; while the latter has 100% TPR (our proposal).